# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v1

@@ F.3 Implementation Details
Copula model. To validate our approach in semi-synthetic datasets, we train two copula models, Clayton and Frank, alongside two nonlinear Weibull proportional-hazards models for the marginal distributions $T_{E}$ and $T_{C}$, respectively. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture. We use a single layer with 32 hidden nodes and the ReLU activation function. For all experiments, we train the models using the Adam optimizer [Kingma and Ba, 2017] with a learning rate of 1e-3, batch normalization enabled, and a batch size of 1024.
Loss function. We estimate the model parameters using maximum likelihood estimation (MLE) under dependent censoring, as described in Foomani et al. [2023]. Given survival data $\mathcal{D}$, let $f_{T\mid X}(\cdot)$ and $F_{T\mid X}(\cdot)$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon, with the event distribution $T_{E}$ and the censoring distribution $T_{C}$. Under independence censoring, the likelihood of the $i$-th instance experiencing the event can then be determined:
$\mathcal{L}(\mathcal{D})=\prod_{i=1}^{N}\left[\text{Pr}(T_{E}=t_{i},T_{C}\geq t_{i}\mid\bm{x}^{(i)})^{\delta^{(i)}}\times\text{Pr}(T_{E}\geq t_{i},T_{C}=t_{i}\mid\bm{x}^{(i)})^{1-\delta^{(i)}}\right]$ (F.3)
However, when $T_{E}$ and $T_{C}$ are no longer independent, we can no longer rely on this clean decomposition of the likelihood. Instead, we introduce a copula $C(u_{1},u_{2})$ to Equation F.3 to adjust for dependent censoring by applying Lemma 2 from Foomani et al. [2023]. Now, as a function of $C(u_{1},u_{2})$ and taking the sum of logs, the (dependent) likelihood has the following form:
$\mathcal{L}(\mathcal{D})=\sum_{i=1}^{N}\delta^{(i)}\;\text{log}\left[f_{T_{E}|\bm{x}^{(i)}}(t_{i}|\bm{x}^{(i)})\right]\;+\;\delta^{(i)}\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T_{E}\rvert\bm{x}^{(i)}}(t_{i}\rvert\bm{x}^{(i)})}_{u_{2}=S_{T_{C}\rvert\bm{x}^{(i)}}(t_{i}\rvert\bm{x}^{(i)})}\right]$ (F.4)
$+\;(1-\delta^{(i)})\;\text{log}\left[f_{T_{C}|\bm{x}^{(i)}}(t_{i}|\bm{x}^{(i)})\right]+(1-\delta^{(i)})\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T_{E}\rvert\bm{x}^{(i)}}(t_{i}\rvert\bm{x}^{(i)})}_{u_{2}=S_{T_{C}\rvert\bm{x}^{(i)}}(t_{i}\rvert\bm{x}^{(i)})}\right]$
@@
Reproducibility. All experiments were implemented and executed in Python 3.9 with PyTorch 1.13.1, NumPy 1.24.3, and Pandas 1.5.3. The Copula-Graphic estimator was implemented using the “compound.Cox” version 3.32 package. Tensor computations were carried out in double precision (fp64). The experiments were conducted on a single workstation equipped with an Intel Core i9-10980XE 3.00GHz CPU, 64GB of RAM, and an NVIDIA GeForce RTX 3090 GPU. Model training was performed on the GPU using CUDA 11.7. All datasets utilized in this study are publicly available. Detailed instructions for reproducing the results are included in the supplementary code.
@@ G.1 Synthetic data
Figures 5-7 demonstrate the bias, i.e., the proximity to the true error, of established baseline metrics and the proposed dependent metrics as a function of the dependence (Kendall’s $\tau$) under a known copula. We plot the mean biases across 10 experiments with different random seeds. Data is generated from a DGP with sample sizes $N=\{1000,5000,10000\}$ and $d=10$. Across all metrics, we generally see larger errors with high dependence ($\tau$). Concerning the small dataset ($N=1,000$), the CI-Dep variant exhibits higher error than Harrell’s CI for both copulas under low to medium dependency ($\tau\leq 0.5)$, however, the CI-Dep improves on both Harrell’s and Uno’s under high dependency ($\tau>0.5)$). The same is the case for the IBS and MAE results, especially under low dependency. In the medium dataset ($N=5,000$), the dependent metrics consistently outperform their independent counterparts across all levels of $\tau$. This trend becomes even more pronounced in the large dataset ($N=10,000$), where the dependent metrics show a significant improvement over the independent ones. These results highlight the increasing impact of dependent censoring as dataset size grows, further demonstrating the limitations of the independent censoring assumption in larger datasets.
@@ H.1 The Kaplan-Meier estimator
Let $(t_{i},\delta_{i}),\;i=1,\ldots,n$, be survival data without covariates, where $t_{i}=\min(T_{i},U_{i})$, $\delta_{i}=I(T_{i}\leq U_{i})$, and $I(\cdot)$ is the censoring bit. We can then consider the survival function $S(t)$ as a function that decreases in steps, with discontinuities occurring only at the observed times of death. In this case, the survival function can be expressed as
$S(t)=\text{Pr}(T\geq t)=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(T=t_{i})\right)\text{Pr}(T\geq t_{i})$ (H.1)
Second, suppose that $T$ and $C$ are independent. Then, we can write
$S(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(T=t_{i},C\geq t_{i})\right)\text{Pr}(T\geq t_{i},C\geq t_{i})$ (H.2)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(\min(T,C)=t_{i},T\leq C)\right)\text{Pr}(\min(T,C)\geq t_{i})$ (H.3)
@@
$S(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{\sum_{\ell=1}^{N}I(t_{\ell}=t_{i},\delta_{\ell}=1)/n}{\sum_{\ell=1}^{N}I(t_{\ell}\geq t_{i})/n}\right)$ (H.4)
@@ H.2 The Copula-Graphic estimator
$P(T>t,C>c)=\varphi^{-1}\big{[}\varphi\{S_{T}(t)\}+\varphi\{S_{C}(c)\}\big{]},$ (H.6)
where $\phi_{\theta}:[0,1]^{d}\rightarrow[0,\infty]$ is a generator function, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3], and $S_{T}(t)=\text{Pr}(T>t)$ and $S_{C}(t)=\text{Pr}(C>c)$ are the marginal survival functions. Let $(t_{i},\delta_{i}),\;i=1,\ldots,n$, be survival data without covariates, where $t_{i}=\min(T_{i},U_{i})$, $\delta_{i}=I(T_{i}\leq U_{i})$, and $I(\cdot)$ is the censoring bit. Assume that $S_{T}(t)$ is a decreasing step function with jumps at event times. Thus, $\delta_{i}=1$ implies $S_{T}(t_{i})\neq S_{T}(t_{i}-dt)$ and $S_{U}(t_{i})=S_{U}(t_{i}-dt)$. Setting $t=u=t_{i}$ in Equation H.6, we have
$\text{Pr}(T\geq t_{i},C\geq t_{i})=\text{Pr}(S_{T}(t_{i}))+\text{Pr}(S_{C}(t_{i}))$ (H.7)
In the left-side of the preceding equation, we estimate $\text{Pr}(T>t_{i},C>t_{i})$ by ($n_{i}-1)/n$, where $n_{i}-1=\sum_{\ell=1}^{n}I(t_{\ell}>t_{i})$ is the number of event-free instances at time $t_{i}$. Then we have
$\phi_{\theta}(\frac{n_{i}-1}{n})=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}.$ (H.8)
We set $t=c=t_{i}-dt$ in Equation H.6 and then estimate $\text{Pr}(T>t_{i}-dt,C>t_{i}-dt)$ by $n_{i}/n$. Then,
$\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{T}(t_{i}-dt)\}+\phi_{\theta}\{S_{C}(t_{i})\},\delta_{i}=1$ (H.9)
@@
$\phi_{\theta}(\frac{n_{i}-1}{n})-\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i}-dt)\},\delta_{i}=1$ (H.10)
We apply the standard constraint that $S_{T}(t_{i}-dt)=1$ when $t_{i}$ is the earliest event time. Under this condition, the solution to the differential equations is
$\phi_{\theta}\{S_{T}(t)\}$ $=\sum_{t_{i}\leq t,\delta_{i}=1}[\phi_{\theta}\{S_{T}(t_{i})\}-\phi_{\delta}\{S_{T}(t_{i}-dt)\}]$ (H.11)
@@
which is equivalent to the CG estimator. Under the independence copula, given by $\phi_{\delta}(t)=-log(t)$, the CG estimator is equivalent to the Kaplan-Meier estimator. Substituting for the Clayton copula [Clayton, 1978], the CG estimator is written as
$S_{T}(t)=\Bigl{[}1+\sum_{t_{i}\leq t,\delta_{i}=1}\Bigl{\{}(\frac{n_{i}-1}{n})^{-\theta}-(\frac{n_{i}}{n})^{-\theta}\Bigr{\}}\Bigr{]}^{-1/\theta}$ (H.13)

### v2

@@ F.3 Implementation Details
Copula model. To validate our approach in semi-synthetic datasets, we train two copula models, Clayton and Frank, alongside two nonlinear Weibull proportional-hazards models for the marginal distributions $E$ and $C$, respectively. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture. We use a single layer with 32 hidden nodes and the ReLU activation function. For all experiments, we train the models for 10000 epochs using the Adam optimizer [Kingma and Ba, 2017] with a learning rate of 1e-3, batch normalization and a batch size of 1024. We use early stopping to terminate the training process if the validation loss does not improve for 100 consecutive epochs.
Loss function. We estimate the model parameters using maximum likelihood estimation (MLE) under dependent censoring, as described in Foomani et al. [2023]. Given survival data $\mathcal{D}$, let $f_{T\mid\bm{X}}(\cdot)$ and $F_{T\mid\bm{X}}(\cdot)$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon, with the event distribution $E$ and the censoring distribution ${C}$. Under independence censoring, the likelihood of the $i$-th instance experiencing the event is:
$\mathcal{L}(\mathcal{D})=\prod_{i=1}^{N}\left[\text{Pr}(E=t_{i},C\geq t_{i}\mid\bm{X}=\bm{x}_{i})^{\delta_{i}}\times\text{Pr}(E\geq t_{i},{C}=t_{i}\mid\bm{X}=\bm{x}_{i})^{1-\delta_{i}}\right]$ (F.3)
However, when $E$ and $C$ are no longer independent, we can no longer rely on this clean decomposition of the likelihood. Instead, we introduce a copula $C(u_{1},u_{2})$ to Equation F.3 to adjust for the dependency between $E$ and $C$ by applying Lemma 2 from Foomani et al. [2023]. Now, as a function of $C(u_{1},u_{2})$ and taking the sum of logs, the (dependent) likelihood has the following form:
$\mathcal{L}(\mathcal{D})=\sum_{i=1}^{N}\delta^{(i)}\;\text{log}\left[f_{{E}\mid\bm{X}}(t_{i}\mid\bm{x}_{i})\right]\;+\;\delta_{i}\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{E\mid\bm{X}}(t_{i}\mid\bm{x}_{i})}_{u_{2}=S_{C\mid\bm{X}}(t_{i}\rvert\bm{x}_{i})}\right]$ (F.4)
$+\;(1-\delta_{i})\;\text{log}\left[f_{{C}\mid\bm{X}}(t_{i}\mid\bm{x}_{i})\right]+(1-\delta_{i})\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{E\rvert\bm{X}}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{{C}\rvert\bm{X}}(t_{i}\rvert\bm{x}_{i})}\right]$
@@
Reproducibility. All experiments were implemented and executed in Python 3.9 with PyTorch 1.13.1, NumPy 1.24.3, and Pandas 1.5.3. The Copula-Graphic estimator was implemented using the compound.Cox package version 3.32. Tensor computations were carried out in double precision (fp64). The experiments were conducted on a single workstation equipped with an Intel Core i9-10980XE 3.00GHz CPU, 64GB of RAM, and an NVIDIA GeForce RTX 3090 GPU. Model training was performed on the GPU using CUDA 11.7. All datasets utilized in this study are publicly available. Detailed instructions for reproducing the results are included in the supplementary code.
@@ G.1 Synthetic data
Figures 5-7 demonstrate the bias, i.e., the proximity to the true error, of established baseline metrics and the proposed dependent metrics as a function of the dependence (Kendall’s $\tau$) under a known copula. We plot the mean biases ($\pm$ SD.) across 100 experiments with different random seeds. Data is generated from a DGP with sample sizes $N=\{1000,5000,10000\}$ and $d=10$. Across all metrics, we generally see larger errors with high dependence ($\tau$). Concerning the small dataset ($N=1,000$), the CI-Dep variant exhibits higher error than Harrell’s CI for both copulas under low to medium dependency ($\tau\leq 0.5)$, however, the CI-Dep improves on both Harrell’s and Uno’s under high dependency ($\tau>0.5)$). The same is the case for the IBS and MAE results, especially under low dependency. In the medium dataset ($N=5,000$), the dependent metrics consistently outperform their independent counterparts across all levels of $\tau$. This trend becomes even more pronounced in the large dataset ($N=10,000$), where the dependent metrics show a significant improvement over the independent ones. These results highlight the increasing impact of dependent censoring as dataset size grows, further demonstrating the limitations of the independent censoring assumption in larger datasets.
@@ H.1 The Kaplan-Meier estimator
Recall that $(t_{i},\delta_{i}),\;i=1,\ldots,n$ represents survival data without covariates, where $t_{i}=\min(e_{i},c_{i})$, $\delta_{i}=\mathbbm{1}[e_{i}\leq c_{i}]$. We can then consider the survival function $S(t)$ as a function that decreases in steps, with discontinuities occurring only at the observed times of death. In this case, the survival function for the event can be expressed as
$S_{E}(t)=\text{Pr}(E\geq t)=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(E=t_{i})\right)\text{Pr}(E\geq t_{i})$ (H.1)
Second, suppose that $E$ and $C$ are independent. Then, we can write
$S_{E}(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(E=t_{i},C\geq t_{i})\right)\text{Pr}(E\geq t_{i},C\geq t_{i})$ (H.2)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(\min(E,C)=t_{i},E\leq C)\right)\text{Pr}(\min(E,C)\geq t_{i})$ (H.3)
@@
$S_{E}(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{\sum_{\ell=1}^{N}I(t_{\ell}=t_{i},\delta_{\ell}=1)/n}{\sum_{\ell=1}^{N}I(t_{\ell}\geq t_{i})/n}\right)$ (H.4)
@@ H.2 The Copula-Graphic estimator
$P(E>t,C>c)=\phi_{\theta}^{-1}\big{[}\phi_{\theta}\{S_{E}(t)\}+\phi_{\theta}\{S_{C}(t)\}\big{]},$ (H.6)
where $\phi_{\theta}:[0,1]^{d}\rightarrow[0,\infty]$ is a generator function, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3], and $S_{E}(t)=\text{Pr}(E>t)$ and $S_{C}(t)=\text{Pr}(C>t)$ are the marginal survival functions. Assume that $S_{E}(t)$ is a decreasing step function with jumps at event times. Thus, $\delta_{i}=1$ implies $S_{E}(t_{i})\neq S_{E}(t_{i}-dt)$ and $S_{C}(t_{i})=S_{C}(t_{i}-dt)$. Setting $t=t_{i}$ in Equation H.6, we have
$\text{Pr}(E\geq t_{i},C\geq t_{i})=\phi_{\theta}\{S_{E}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}$ (H.7)
In the left-side of the preceding equation, we estimate $\text{Pr}(E>t_{i},C>t_{i})$ by ($n_{i}-1)/n$, where $n_{i}-1=\sum_{\ell=1}^{n}\mathbbm{1}[t_{\ell}>t_{i}]$ is the number of event-free instances at time $t_{i}$. Then we have
$\phi_{\theta}(\frac{n_{i}-1}{n})=\phi_{\theta}\{S_{E}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}.$ (H.8)
We set $t=c=t_{i}-dt$ in Equation H.6 and then estimate $\text{Pr}(E>t_{i}-dt,C>t_{i}-dt)$ by $n_{i}/n$. Then,
$\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{E}(t_{i}-dt)\}+\phi_{\theta}\{S_{C}(t_{i})\},\delta_{i}=1$ (H.9)
@@
$\phi_{\theta}(\frac{n_{i}-1}{n})-\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{E}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i}-dt)\},\delta_{i}=1$ (H.10)
We apply the standard constraint that $S_{E}(t_{i}-dt)=1$ when $t_{i}$ is the earliest event time. Under this condition, the solution to the differential equations is
$\phi_{\theta}\{S_{E}(t)\}$ $=\sum_{t_{i}\leq t,\delta_{i}=1}[\phi_{\theta}\{S_{E}(t_{i})\}-\phi_{\delta}\{S_{E}(t_{i}-dt)\}]$ (H.11)
@@
which is equivalent to the CG estimator. Under the independence copula, given by $\phi_{\delta}(t)=-\log(t)$, the CG estimator is equivalent to the Kaplan-Meier estimator. Substituting for the Clayton copula [Clayton, 1978], the CG estimator is written as
$S_{E}(t)=\Bigl{[}1+\sum_{t_{i}\leq t,\delta_{i}=1}\Bigl{\{}(\frac{n_{i}-1}{n})^{-\theta}-(\frac{n_{i}}{n})^{-\theta}\Bigr{\}}\Bigr{]}^{-1/\theta}$ (H.13)